a 4% salt solution is mixed with an 8% salt solution. How many ounces of each solution are needed to obtain 400 oz. of a 5% solution? I need to solve using linear equations.
Answer by Alan3354(69443)

a 4% salt solution is mixed with an 8% salt solution. How many ounces of each solution are needed to obtain 400 oz. of a 5% solution? I need to solve using linear equations.
--------------
Do it like this one:
if a 14% acid solution is mixed with a 19% acid solution to produce 45 gallons of a 15% acid solution how much (in gallons) of each solution should be used?
--------------
f = 14%
n = 19%
----
f + n = 45 (total gallons)
14f + 19n = 15*45 (total acid)
-----
n = 45 - f
Sub for n
14f + 19*(45-f) = 675
14f + 855 - 19f = 675
-5f = -180
f = 36 gallons of 14%
n = 9 gallons of 19%